Circular No. 7160

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 990510 M. Dadina, L. Di Ciolo, and A. Coletta, BeppoSAX, Rome; P. Soffitta

and M. Feroci, Istituto Astrofisica Spaziale, CNR, Rome; J. Heise and E. Kuulkers, Space Research Organization of the Netherlands (SRON), Utrecht; J. M. Muller, BeppoSAX, Rome, and SRON; and F. Frontera and L. Amati, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna, write: “The BeppoSAX Gamma Ray Burst Monitor observed a γ-ray burst starting on May 10.36743 UT. A preliminary analysis reveals a duration of ∼ 80 s and a multipeak structure, with a peak intensity of 3100 counts/s in the energy band 40–700 keV. The burst was also detected in Wide Field Camera unit 2 with comparable duration, a peak intensity of ∼ 4.3 Crab and an average intensity of 0.7 Crab (2–28 keV). The position of the x-ray counterpart is α = 13^{h}38^{m}06^{s}, δ = −80^{o}29′.5 (equinox 2000.0), with an error radius of 3′. A follow-up observation with BeppoSAX narrowfield instruments is in progress.”

SUPERNOVA 1999ca IN NGC 3120 F. Patat, European Southern Observatory (ESO); J. Maza, Univer-

sidad de Chile; S. Benetti, Telescopio Nazionale Galileo, Centro Galileo Galilei, Santa Cruz de La Palma; and E. Cappellaro, Padova Observatory, write: “We have observed SN 1999ca (cf. IAUC 7158) with the ESO New Technology Telescope (+ EMMI; range 340–1000 nm, resolution 1.0 nm) on May 7.05 UT at La Silla. A full reduction of the CCD spectrogram shows that this is a type-II supernova around one month after maximum light. Hα, Hβ, Hγ, and Hδ, superimposed on a continuum that peaks around 550 nm, are clearly visible. The Hα line is quite asymmetric, with possible unresolved structures, and its P-Cyg absorption is very weak. Besides the Balmer series, the most prominent features are identified as Ca II, Ba II, Fe I, Na I, and Ca II. The expansion velocities, deduced from the absorption minima, are 6800, 7600, and 6400 km/s for Ca II H and K, Na I D, and Hα, respectively. Two unresolved absorptions of Na I D are clearly visible; one is at the parent-galaxy rest-frame velocity (2791 km/s; Theureau et al. 1998, A.Ap. Suppl. 130, 333), while the other absorption most likely originates in the Milky Way. The total equivalent widths are 0.27 and 0.13 nm for the two Na features, respectively.”

1999 May 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green